OpenStudy (anonymous):

Solve. s 18 - = - 39 26 s =

OpenStudy (igreen):

Cross multiply. s * 26 = 39 * 18

OpenStudy (anonymous):

39*18=702

OpenStudy (anonymous):

702*26=18,252

OpenStudy (anonymous):

like that

OpenStudy (igreen):

No..

OpenStudy (igreen):

We don't multiply 26..we divide it

OpenStudy (anonymous):

27

OpenStudy (anonymous):

702 divided by 26 = 27

OpenStudy (anonymous):

is it right

OpenStudy (igreen):

Yes, correct.